OpenStudy (anonymous):

use quadratic formula to solve x^2-5x+6=0

OpenStudy (anonymous):

\[x^2-5x+6= (-(-5)\pm \sqrt{(-5)^2-4(1)(6)})/2(1)\] \[(5\pm \sqrt{25-24})/2\] \[(5\pm \sqrt{1})/2\] \[5 + 1= 6/2 = 3 and 5-1=4/2= 2\] Solutions are x = 3; x = 2